# F-Fidelity: A Robust Framework for Faithfulness Evaluation of Explainable AI

### v1

@@ Appendix C Detailed Experiment Setup
We explore two architectures, Resnet and ViT for the image classification tasks. We use Resnet-9  [^3] https://jovian.com/tessdja/resnet-practice-cifar100-resnet and a 6-layer ViT as backbones for both  Cifar-100  and  Tiny-Imagenet . In the ResNet architecture, the hidden dimensions are set to $[64,128,128,256,512,512,512,512]$. For the ViT model, we use a patch size of 4, with a patch embedding dimension of 128. The backbone consists of 6 attention layers, each with 8 heads for multi-head attention. The final output hidden dimension of the ViT encoder is 256. In the training stage, we set the learning rate and weight decay to 1E-4 for Resnet and set the learning rate to 1E-3 and weight decay to 1E-4 for ViT. We use Adam as the optimizer and the training epochs are 100 for Resnet and 200 for ViT. During fine-tuning, we use the same hyperparameters. We report our model accuracy in the Table 4.
Table 4: The accuracy of the models to be explained in the image classification task.
@@
Dataset origin retrain fine-tune origin retrain fine-tune
Cifar-100 0.7323 0.7001 0.7310 0.5025 0.5071 0.4691
Tiny-Imagenet 0.5585 0.5600 0.5476 0.4255 0.4262 0.3825
Table 5: The accuracy of the models to be explained in the time series classification task.
LSTM Transformer
Dataset origin retrain fine-tune origin retrain fine-tune
PAM 0.8773 0.8446 0.8886 0.8783 0.8502 0.8689
Boiler 0.7975 0.7750 0.7425 0.8425 0.7575 0.8325
For the time series classification task, we follow the TimeX (Queen et al., 2023) to use a simple LSTM for both Boiler and PAM datasets. The simple LSTM model contains 3 bidirectional LSTM layers with 128 hidden embedding sizes. We use AdamW as the optimizer with a learning rate of 1E-3 and weight decay is 1E-2. Similarly, we provide the model accuracy in the Table 5.
Appendix D Extra Experimental Study
D.1 Ablation Study
To evaluate the effectiveness of our proposed F-Fidelity framework, we conduct ablation studies comparing three evaluation setups: the original Fidelity metric, Fidelity with fine-tuning (Fidelity+Fine-tune), and F-Fidelity. We utilize the CIFAR dataset for these experiments, employing a ResNet architecture as the backbone for the classifier. The explainers are based on SG-SQ.
@@ D.1 Ablation Study
D.2 Hyperparameter Sensitivity Study
In this part, we assess the robustness of F-Fidelity to hyperparameter choices. We select a subset of 400 samples with an explanation size ratio of 0.2 from the colored-MNIST dataset (Arjovsky et al., 2019). we have three key hyperparameters: the number of sampling iterations $T$, the ratio $\alpha^{+}/\alpha^{-}$, and $\beta$. For each hyperparameter, we explored a range of values to understand their impact on the performance of F-Fidelity. We evaluated the method’s performance using three Macro Spearman correlations, with additional detailed results on $RFid^{+}$, $RFid^{-}$, and Micro Spearman correlations provided in Appendix E.4.
(a) Ablation study on $T$
(b) Ablation study on $\beta$
(c) Ablation study on $\alpha$
Figure 3: Ablation study on sampling number $T$, $\beta$, and $\alpha=\alpha^{+}=\alpha^{-}$. We report the Macro Spearman rank correlations.
D.3 Fidelity and Explanation Size
As mentioned in Section D.2, We conduct experiments on three settings of $\alpha^{+}$ and five settings of $\beta$ in the evaluation stage. As Figure 4 shows, in different settings, the ground truth size can be observed from the flat area in $RFid^{+}$. According to the analysis in Section 5, there are two key points in the x-axis, the explanation size $\gamma$ and the point where the deleting attributions are equal to the explanation size, $\beta\/\alpha^{+}$. Between the two key points, there exists a flat area, which is the best comparison area. Notably, our experiments are based on the assumption that all the explanation parts share the same important weight. In the real world, there is a gap between the ideal results and experiment results, such as the smooth increase and decrease in Figure  4.

### v2

@@ Appendix C Detailed Experiment Setup
We explore two architectures, ResNet and ViT for the image classification tasks. We use ResNet-9  [^3] https://jovian.com/tessdja/resnet-practice-cifar100-resnet and a 6-layer ViT as backbones for both  CIFAR-100  and  Tiny-Imagenet . In the ResNet architecture, the hidden dimensions are set to $[64,128,128,256,512,512,512,512]$. For the ViT model, we use a patch size of 4, with a patch embedding dimension of 128. The backbone consists of 6 attention layers, each with 8 heads for multi-head attention. The final output hidden dimension of the ViT encoder is 256. In the training stage, we set the learning rate and weight decay to 1E-4 for ResNet and set the learning rate to 1E-3 and weight decay to 1E-4 for ViT. We use Adam as the optimizer and the training epochs are 100 for ResNet and 200 for ViT. During fine-tuning, we use the same hyperparameters.
For the time series classification task, we follow the TimeX (Queen et al., 2023) to use a simple LSTM for both Boiler and PAM datasets. The simple LSTM model contains 3 bidirectional LSTM layers with 128 hidden embedding sizes. We use AdamW as the optimizer with a learning rate of 1E-3 and weight decay is 1E-2.
For the natural language task, we use LSTM (Hochreiter & Schmidhuber, 1997) and Transformer as the backbone. the LSTM has one hidden layer with the dimension set to 128. In Transformer, the hidden dimension is set to 512. The number of Transformer layers for SST2 and BoolQ are 2 and 4. The head of the Transformer layers for SST2 and BoolQ are 4 and 8. For all datasets and architectures, we use Adam optimizer (Kingma & Ba, 2015) with default learning rate 1E-4, training epochs 100.
Appendix D Extra Experimental Study
D.1 Experiments with Other Backbones.
ViT as Backbone for Image Classification. To further validate the robustness of our F-Fidelity framework across different model architectures, we conducted additional experiments using ViT (Dosovitskiy et al., 2020) as the backbone for image classification on both CIFAR-100 and TinyImageNet datasets. We utilize the SG-SQ (Smilkov et al., 2017) to generate explanations. For the settings, we follow our main experimental setup. As shown in Table 5, F-Fidelity consistently achieves best correlations across all metrics for both datasets, outperforming other methods. This strong performance on ViT models further emphasizes the versatility and effectiveness of F-Fidelity in evaluating explanations across different deep learning architectures.
Correlation CIFAR-100 Tiny-Imagenet
CNN as Backbone for Time Series Classification. To verify the effectiveness of F-Fidelity on Time Series, we conduct experiments with CNN as the classifier by following the setting in Section  5.2. As Table 6 shows, F-Fidelity consistently outperforms baselines across all metrics including macro and micro correlations. Similarly, we observe the same phenomenon as shown in Table 3, the Fidelity and ROAR suffer OOD problem. Under some sparsity, the correlation is not available for the same rank while F-Fidelity and R-Fidelity perform well. These results on CNN models highlight the versatility and effectiveness of F-Fidelity in assessing explanations across various deep learning architectures.
Table 6: Spearman ranks correlations and ranks on time series datasets with CNN as the classifier. The best performance is marked as bold. The “-” means the correlation can’t be obtained because the different explanations have the same rank.
Correlation PAM Boiler
Macro Corr. MoRF vs GT $\downarrow$ 0.78$\pm$0.10 -0.90$\pm$0.06 -1.$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
LeRF vs GT $\uparrow$ 0.96$\pm$0.14 -0.71$\pm$0.10 1.00$\pm$0.00 1.00$\pm$0.00 0.91$\pm$0.07 0.86$\pm$0.07 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 0.68$\pm$0.18 0.61$\pm$0.11 -1.$0$1.$0$0.91$\pm$0.07 -0.86$\pm$0.07 -1.$0$1.00$\pm$0.00
Micro Corr. MoRF vs GT $\downarrow$ -0.02$\pm$0.62 -0.39$\pm$0.66 -1.$0$1.$0$0.94$\pm$0.13 -0.92$\pm$0.18
LeRF vs GT $\uparrow$ 0.93$\pm$0.15 0.68$\pm$0.53 1.00$\pm$0.01 1.$0$ 0.93$\pm$0.11 0.93$\pm$0.11
MoRF vs LeRF $\downarrow$ -0.03$\pm$0.58 -0.12$\pm$0.61 -1.$0$1.$0$0.93$\pm$0.12 -0.94$\pm$0.11
Micro Rank MoRF vs GT $\downarrow$ 3.00$\pm$1.26 2.00$\pm$1.08 1.00$\pm$0.00 1.11$\pm$0.45 3.16$\pm$0.87 2.63$\pm$0.87 1.11$\pm$0.31 1.32$\pm$0.46
LeRF vs GT $\downarrow$ 2.68$\pm$1.22 2.84$\pm$1.42 1.00$\pm$0.00 1.00$\pm$0.00 3.11$\pm$0.64 3.53$\pm$0.50 1.47$\pm$0.60 1.21$\pm$0.41
MoRF vs LeRF $\downarrow$ 3.68$\pm$0.46 2.79$\pm$1.15 1.00$\pm$0.00 1.11$\pm$0.45 3.21$\pm$0.41 3.05$\pm$0.22 1.32$\pm$0.46 1.16$\pm$0.36
Transformer as Backbone for Natural Language Classification. To further validate our findings with modern architectures, we conducte experiments using Transformer models on both SST2 and BoolQ datasets. As shown in Table 7, the results demonstrate interesting patterns. On SST2, while all methods perform well, F-Fidelity shows slight advantages, particularly in micro correlations and rankings. For BoolQ, all methods achieve nearly perfect correlations, suggesting that Transformer models on this dataset exhibit strong robustness to perturbations, making the differences between evaluation methods less pronounced. These results align with our earlier observations about the inherent robustness of NLP models, while still demonstrating the reliability and slight advantages of F-Fidelity in more challenging scenarios.
Table 7: Spearman rank correlation and rank results with Transformer model on SST2 and BoolQ datasets.
Correlation SST2 BoolQ
Fidelity R-Fidelity F-Fidelity Fidelity R-Fidelity F-Fidelity
Macro Corr. MoRF vs GT $\downarrow$ -$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
Micro Corr. MoRF vs GT $\downarrow$ -0.93$\pm$0.07 -0.97$\pm$0.05 -1.00$\pm$0.01 -1.00$\pm$0.01 -1.00$\pm$0.01 -1.00$\pm$0.00
LeRF vs GT $\uparrow$ 0.98$\pm$0.04 0.99$\pm$0.02 0.99$\pm$0.03 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -0.92$\pm$0.09 -0.97$\pm$0.05 -0.99$\pm$0.02 -1.$0$1.$0$1.00$\pm$0.00
Micro Rank MoRF vs GT $\downarrow$ 2.05$\pm$0.89 1.32$\pm$0.46 1.05$\pm$0.22 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
LeRF vs GT $\downarrow$ 1.16$\pm$0.49 1.26$\pm$0.55 1.16$\pm$0.36 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 2.05$\pm$0.97 1.32$\pm$0.46 1.11$\pm$0.45 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
D.2 Relationship Between Model Robustness with F-Fidelity Performance
To understand how model robustness influences F-Fidelity’s effectiveness, we conduct a systematic evaluation across three domains: computer vision (CIFAR-100), time series (Boiler), and NLP (SST2). For each domain, we test multiple architectures - ResNet and Transformer for CIFAR-100, LSTM and CNN for Boiler, and LSTM and Transformer for SST2. We measure model robustness through classification accuracy under different perturbation levels (0%, 5%, and 10%), where perturbations are implemented through domain-specific masking operations: zeroing pixels for images, masking feature values for time series, and replacing token with zeros for text. For each model-architecture pair, we compare the performance between the original model and its fine-tuned version under these perturbation conditions.
Table 8: Classification accuracy (%) comparison between original and fine-tuned models under different perturbation ratios.
Dataset Architecture Model 0% 5% 10%
CIFAR-100 ResNet Original 73.23 56.70 41.81
Fine-tuned 73.10 73.12 72.97
ViT Original 50.25 46.19 39.87
Fine-tuned 46.91 47.05 47.30
Boiler LSTM Original 80.14 67.78 61.53
Fine-tuned 77.22 72.64 73.19
CNN Original 82.22 68.89 57.92
Fine-tuned 86.39 74.03 63.89
SST2 LSTM Original 82.45 80.28 79.82
Fine-tuned 82.68 80.16 79.24
Transformer Original 80.85 77.29 73.28
Fine-tuned 80.73 81.08 78.90
Table 8 reveals distinct robustness patterns across different domains and architectures. In computer vision, models exhibit high sensitivity to perturbations - the original ResNet on CIFAR-100 shows dramatic accuracy degradation under even small perturbations, while its fine-tuned counterpart maintains stable performance. Time series models demonstrate moderate sensitivity, with original models showing notable but less severe accuracy drops under perturbations, which fine-tuning helps to stabilize. In contrast, NLP models display inherent robustness, maintaining relatively stable performance even under perturbations regardless of fine-tuning.
These robustness patterns directly correlate with F-Fidelity’s effectiveness. The performance improvements provided by F-Fidelity are most pronounced in domains where models show high sensitivity to perturbations (computer vision and time series). This is because traditional fidelity metrics in these domains suffer from unreliable evaluations due to OOD inputs, which our fine-tuning strategy effectively mitigates. However, in NLP tasks where models possess natural robustness to perturbations, the OOD problem is less severe, consequently diminishing the relative advantages of F-Fidelity. This relationship provides practical guidance for choosing evaluation metrics - F-Fidelity should be preferred in domains where models show high sensitivity to perturbations, while simpler metrics may suffice for naturally robust domains.
D.3 Analysis of Decision Boundary Preservation under Random Masking Fine-Tuning
This section addresses whether fine-tuning the classification model with random masking (up to fraction $\beta$ of input elements) significantly changes the model’s decision boundaries. We examine this by looking at both global classification patterns and local decision characteristics. We consider two aspects of the model’s decision boundaries as follows.
  • Global Decision Boundary: The global decision boundary refers to the overall separation between different classes in the input space. It captures the main structure of the data distribution and represents the classifier’s ability to distinguish between classes on a broad scale.
  • Local Decision Boundary: The local decision boundary characterizes how the model responds to small perturbations around individual data points.
(a) Classification accuracy
(b) Prediction agreement with original prediction
Figure 1: Effects of random masking fine-tuning on model behavior. Higher fine-tuning $\beta$ values improve robustness to perturbations but reduce agreement with the original model’s predictions.
Our random masking fine-tuning aims to maintain the model’s global classification behavior while improving its stability to small input changes. The random masking process applies uniform noise across all inputs, regardless of their class. This helps the model learn to handle missing information while keeping its main classification patterns intact. Random masking fine-tuning maintains model performance through several key principles. First, the random masking operation applies a uniform noise distribution across the input space. Since this noise is class-agnostic and uniformly distributed, it preserves the statistical properties of the underlying class distributions. Second, random masking acts as a form of dropout regularization (Srivastava et al., 2014), which has been shown to improve model generalization while maintaining core decision boundaries. Third, the introduction of masked inputs during training encourages the model to develop smoother decision boundaries in local neighborhoods, as demonstrated by Bishop (1995) in the context of noise injection.
To empirically evaluate random masking fine-tuning, we conduct experiments with ResNet on CIFAR-100. We examine both global decision boundary preservation and local boundary characteristics through systematic perturbation analysis. For global boundary preservation, we compare classification accuracy between the original and fine-tuned models on clean test data, using 0% masking as the baseline reference point. We also measure prediction agreement between these models to assess consistency in their decision-making. To analyze local boundary characteristics, we apply incremental perturbations during testing by randomly masking input elements, with perturbation sizes ranging from 0.05 to 0.50. At each perturbation level, we track both classification accuracy and prediction agreement with the original model. Classification accuracy reveals the model’s robustness to perturbations, while prediction agreement quantifies the consistency of decision boundaries under noise. The masking ratio $\beta$ during fine-tuning controls the balance between preserving global boundaries and smoothing local ones.
Our experimental results demonstrate that random masking fine-tuning effectively balances global decision boundary preservation with local boundary smoothing. The fine-tuned models maintain comparable baseline accuracy to the original model while showing significantly improved robustness to perturbations. On clean test data, the accuracy difference between original and fine-tuned models remains small, indicating preserved global classification behavior. The high prediction agreement between original and fine-tuned models on clean data further confirms the preservation of global decision patterns. Meanwhile, the smoother local decision boundaries manifest through sustained performance under perturbations, with fine-tuned models showing notably accuracy change compared to the original model.
The masking ratio $\beta$ plays a crucial role in model behavior under perturbations, as evidenced in Figure 1. In terms of accuracy (Figure 1(a)), models fine-tuned with larger $\beta$ values demonstrate remarkable resilience, maintaining accuracy above 60% even under 50% input perturbation, while the original model ($\beta=0$) drops below 40% accuracy with just 10% perturbation. The prediction agreement results (Figure 1(b)) show a trade-off. Higher $\beta$ values lead to better accuracy under perturbations and result in lower agreement with the original model’s predictions. This suggests that stronger fine-tuning with larger $\beta$ values causes the model to learn more robust but slightly different decision boundaries compared to the original model. Models with moderate $\beta$ values (like 0.10) maintain a better balance between perturbation resistance and consistency with the original model’s behavior.
D.4 Preservation of Explanation After Fine-tuning
To further validate that our fine-tuning strategy preserves meaningful explanations, we conduct a qualitative analysis on Tiny-Imagenet. We visualize and compare the attribution maps generated for both the original and fine-tuned models using GradCAM. As shown in Table 9, the attribution maps from both models highlight similar regions of importance, with only minor variations in the highlight intensity. This visual consistency suggests that while our fine-tuning process improves model robustness to perturbations, it maintains the model’s original attention to meaningful features, thereby preserving the validity of the generated explanations.
Table 9: The explanation visualization comparison between original and fine-tuned model.
@@ D.4 Preservation of Explanation After Fine-tuning
Input Original Fine-tuned Original Fine-tuned
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
D.5 Evaluation with LayerCAM as Alternative Explanation Method
To further validate the robustness of our method across different explanation techniques, we conduct additional experiments using LayerCAM (Jiang et al., 2021), an alternative explanation method to GradCAM. Following the same experimental setup as in Table 1, we evaluate the performance using ResNet architecture on the CIFAR-100 dataset. As shown in Table 10, F-Fidelity maintains its superior performance with LayerCAM, exhibiting consistent patterns with our GradCAM results. This consistency across different explanation methods reinforces the generalizability of our evaluation framework and its ability to assess various types of explanation techniques reliably.
Table 10: Spearman rank correlation and rank results on CIFAR-100 dataset with ResNet explained by LayerCAM.
Correlation Fidelity ROAR R-Fidelity F-Fidelity
Macro Corr. MoRF vs GT $\downarrow$ 0.77$\pm$0.06 0.98$\pm$0.03 $0$0.26$\pm$0.00
LeRF vs GT $\uparrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 0.77$\pm$0.06 0.98$\pm$0.03 $0$0.26$\pm$0.00
Micro Corr. MoRF vs GT $\downarrow$ 0.53$\pm$0.40 0.81$\pm$0.20 -0.02$\pm$0.22 -0.50$\pm$0.61
LeRF vs GT $\uparrow$ 0.99$\pm$0.02 0.99$\pm$0.04 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 0.53$\pm$0.41 0.81$\pm$0.20 -0.02$\pm$0.22 -0.50$\pm$0.61
Micro Rank MoRF vs GT $\downarrow$ 2.95$\pm$0.23 3.95$\pm$0.23 1.89$\pm$0.66 1.21$\pm$0.42
LeRF vs GT $\downarrow$ 2.53$\pm$1.39 1.95$\pm$1.18 1.05$\pm$0.23 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 2.89$\pm$0.32 3.95$\pm$0.23 1.89$\pm$0.56 1.21$\pm$0.42
D.6 Case Study
To provide the intuitional verification of our method, we conduct a case study with an image from the Tiny-Imagenet dataset. We first use GradCAM to obtain the explanation. To get various degraded explanations, we follow the previous setting to add noise, which is a list of $[0.2,0.4,0.6,0.8,1.0]$, to the explanation to obtain the ground truth ranking of explanations. We compare baselines and F-Fidelity and show the results in Figure 2. We observe that the Fidelity and ROAR methods fail to differentiate explanations with noise levels of 0.2 and 0.4 when the sparsity is below certain thresholds (e.g., 25%). However, the R-Fidelity and F-Fidelity have a good performance in such cases and F-Fidelity has a better performance than R-Fidelity from the micro correlation results.
Figure 2: Case study of a sample in Tiny-Imagenet. We use GradCAM to obtain the explanation and get degraded explanations by adding random noise to the explanation. The noise level is a list of $[0.2,0.4,0.6,0.8,1.0]$. We visualize the Fidelity score and Correlation on different sparsity.
D.7 Ablation Study
To evaluate the effectiveness of our proposed F-Fidelity framework, we conduct ablation studies comparing three evaluation setups: the original Fidelity metric, Fidelity with fine-tuning (Fidelity+Fine-tune), and F-Fidelity. We utilize the CIFAR-100 dataset for these experiments, employing a ResNet architecture as the backbone for the classifier. The explainers are based on SG-SQ.
@@ D.7 Ablation Study
D.8 Emprical Verification of determining the Explanation Size
Following the terminology of Section 4, the pixels can be divided into two tiers: those belonging to the digit (forming the explanation) and those in the background. The explanation size is thus $c_{1}=\gamma td$. According to Theorem 1, we expect $FFid^{+}$ to increase for $s<c_{1}$ and decrease for $s>\max\left(\frac{\beta}{\alpha^{+}}td,\gamma td\right)$.
To verify this behavior, we evaluate $FFid^{+}$ across three settings of $\alpha^{+}$ and five settings of $\beta$. As shown in Figure 4, the ground truth explanation size can be identified from the flat area in $RFid^{+}$ curves. Two critical points emerge on the x-axis: the explanation size $\gamma$ and the point where the removal attributions equal the explanation size ($\beta/\alpha^{+}$). The region between these points forms a plateau that enables reliable comparison of explanations. When $\frac{\beta}{\alpha^{+}}td<c_{1}$, the point where $FFid^{+}$ changes direction corresponds to the explanation size $\gamma$, allowing us to recover the true explanation size by evaluating $FFid^{+}$ over different $\beta$ values.
While our theoretical analysis assumes uniform importance weights across explanation components, real-world applications show some deviation from this idealized behavior, manifesting as smooth transitions rather than sharp changes in the $FFid^{+}$ curves. Nevertheless, the empirical results validate the theoretical predictions and demonstrate the utility of $FFid$ metrics in determining explanation sizes.